# Large Language Models for Data Annotation: A Survey

### v2

@@ Abstract
Data annotation generally refers to the labeling or generating of raw data with relevant information, which could be used for improving the efficacy of machine learning models. The process, however, is labor-intensive and costly. The emergence of advanced Large Language Models (LLMs), exemplified by GPT-4, presents an unprecedented opportunity to automate the complicated process of data annotation. While existing surveys have extensively covered LLM architecture, training, and general applications, we uniquely focus on their specific utility for data annotation. This survey contributes to three core aspects: LLM-Based Annotation Generation, LLM-Generated Annotations Assessment, and LLM-Generated Annotations Utilization. Furthermore, this survey includes an in-depth taxonomy of data types that LLMs can annotate, a comprehensive review of learning strategies for models utilizing LLM-generated annotations, and a detailed discussion of the primary challenges and limitations associated with using LLMs for data annotation. Serving as a key guide, this survey aims to assist researchers and practitioners in exploring the potential of the latest LLMs for data annotation, thereby fostering future advancements in this critical field.
@@ 1 Introduction
In the complex realm of machine learning and natural language processing (NLP), data annotation stands out as a critical yet challenging task, extending beyond simple label attachment to encompass a diverse array of fundamental or auxiliary information. This detailed process typically involves ❶ categorizing raw data with class or task labels for basic classification, ❷ adding intermediate labels for contextual depth Yu et al. (2022), ❸ assigning confidence scores to assess annotation reliability Lin et al. (2022), ❹ applying alignment or preference labels to tailor outputs to specific criteria or user needs, ❺ annotating entity relationships to understand how entities within a dataset interact with each other Wadhwa et al. (2023), ❻ marking semantic roles to define the underlying roles that entities play in a sentence Larionov et al. (2019), or ❼ tagging temporal sequences to capture the order of events or actions Yu et al. (2023).
Despite its wide applications, data annotation poses significant challenges for current machine learning models due to the complexity, subjectivity, and diversity of data. This process requires domain expertise and is resource-intensive, particularly when manually labeling large datasets. Advanced LLMs such as GPT-4 OpenAI (2023), Gemini Team et al. (2023), and LLaMA-2 Touvron et al. (2023b) offer a promising opportunity to revolutionize data annotation. LLMs serve as more than just tools but play a crucial role in improving the effectiveness and precision of data annotation. Their ability to automate annotation tasks (Zhang et al., 2022), ensure consistency across large volumes of data (Hou et al., 2023), and adapt through fine-tuning or prompting for specific domains (Song et al., 2023), significantly mitigates the challenges encountered with traditional annotation methods, setting a new standard for what is achievable in the realm of NLP. This survey delves into the nuances of using LLMs for data annotation, exploring methodologies, utilizing strategies, and associated challenges in this transformative approach. Through this exploration, we aim to shed light on the motivations behind embracing LLMs as catalysts for redefining the landscape of data annotation in machine learning and NLP. We explore the utilization of LLMs for data annotation in this survey, making four main contributions:
  • LLM-Based Annotation Generation: We dive into the process of generating annotations for various data types, including instruction & response, rationale, pairwise feedback, textual feedback, and other domain-specific data. Additionally, we discuss the criteria (e.g., diversity and quality) in the annotation process.
@@
Differences from Other LLM-related Surveys. While existing surveys in the NLP domain extensively cover architectural nuances Zhao et al. (2023a), training methodologies Liu et al. (2023d), and evaluation protocols Chang et al. (2023) associated with LLMs, their main focus lies on the capabilities of models for specific end tasks such as machine translation Min et al. (2021), alignment Wang et al. (2023f), code generation Zan et al. (2023), and medical analysis Thirunavukarasu et al. (2023). In contrast, this survey distinguishes itself by focusing primarily on the application of these potent next-generation LLMs to the intricate realm of data annotation, a domain that is crucial yet underexplored.
@@ 2 Preliminaries
In this section, we delve into our approach to the annotation process. We introduce two core models: an annotator model, denoted as $\mathcal{A}$, which maps input data to annotations, and a task learner, represented as $\mathcal{L}$, that utilizes or learns from these annotated data to accomplish specific tasks. Our primary focus is on utilizing advanced LLMs like GPT-4 (OpenAI, 2023) and LLaMA (Touvron et al., 2023a) as annotators ($\mathcal{A}$), while the task learner ($\mathcal{L}$) can be another large model Chiang et al. (2023a) or a less complex one such as BERT (Devlin et al., 2018), which utilizes these annotated data to perform designated tasks. LLM-generated annotations encompass categorical labels and enhance raw data points with a comprehensive array of auxiliary signals. These annotations, including confidence scores, contextual details, and other metadata, extend beyond traditional categorical labels.
@@ 3 LLM-Based Annotation Generation
The emergence of LLMs has sparked significant interest in their capacity for high-quality, context-sensitive data annotation. This section discusses various kinds of annotations produced via LLMs.
@@ 3.1 Instruction & Response
Instruction and response are the two fundamental components that constitute a dataset for LLM fine-tuning and in-context learning (ICL). Previous NLP datasets Li et al. (2017); Wang et al. (2018); Ouyang et al. (2022) mainly rely on human annotators to construct. Recently, with the advent of LLMs, automatic and generative methods Meng et al. (2022); Ye et al. (2022a, b); Wang et al. (2024d) have gained more focus in data annotation.
@@
Instruction Diversity. The diversity of instruction has been proven crucial for LLM learning Li et al. (2023e); Song et al. (2024b, a). Recent studies have explored various methods to diversify and augment instructions in the original datasets. For example, Yoo et al. (2021) enhance data diversity by mixing two different samples to create a new one. Wang et al. (2022b) use a few manually-written seed instructions and iteratively augment them with a generate-then-filter pipeline. Additionally, Meng et al. (2023); Wang et al. (2024c) train an instruction generation model in the original dataset to augment the diversity of instruction. Gupta et al. (2023) employ a multi-step prompting method to first generate task descriptions, which are then used as instance seeds to guide LLMs in instruction generation. To obtain informative and diverse examples, Wang et al. (2023c) propose an explain-then-generate pipeline with LLMs for iterative data synthesis. Besides, Li et al. (2023a) paraphrase the given sample multiple times to help LLMs understand them from different perspectives. Köksal et al. (2024) suggest a clustering-based data selection method to ensure diversity in the initial seed data for augmentation. Recently, Yu et al. (2024) introduce AttrPrompt as an effective way to balance diversity and cost in LLM-based data annotation.
@@
Response Quality. High-quality responses are essential for effective fine-tuning and ICL Luo et al. (2024). To improve the quality of the generated response, Zhang and Yang (2023a) frame the response generation as reading comprehension tasks and create detailed prompts for LLMs. Huang et al. (2023) adopt self-consistency Wang et al. (2022b) in response generation, selecting from the candidate response with the highest confidence score. Furthermore, Yang et al. (2024b) propose self-distill and augment the instruction tuning dataset by rewriting the original responses. Pang et al. (2024b) conduct social simulations to ensure high-quality, human-valued responses from LLMs. Moreover, Liu et al. (2024) introduce a multi-step prompting including question analysis, answer guidance and safe answer production in their response generation pipeline. Guo et al. (2024a) enhance the LLMs outputs’ quality by implementing retrieval-augmented ICL and providing LLMs with relevant documents. To ensure LLMs provide responses aligned with human values, Sun et al. (2024b) and Wang et al. (2024a) conduct principle-driven prompting, guiding LLMs with well-crafted and detailed principles.
@@ 3.2 Rationale
Rationale Structure. Following Kojima et al. (2022), there is a notable interest in abstracting the reasoning process of LLMs into diverse structures and format, including trees Hao et al. (2023); Yao et al. (2024), graphs Besta et al. (2024); Yao et al. (2023), tables Wang et al. (2024e), programs Chen et al. (2023e), recursion Qi et al. (2023), and concepts Tan et al. (2023).
Rationale Quality. To produce high-quality and fine-grained rationale, diverse methodologies have been employed. Wang et al. (2022a) prompt frozen LLMs to produce choice-specific rationales to elucidate each choice in a sample. Wang et al. (2023b) employ contrastive decoding to foster more plausible rationales, taking into account gold-standard answers. Liu et al. (2023a) curate meticulously designed prompts to derive high-quality rationales from GPT-4 and construct a logical CoT instruction tuning dataset. For attaining fine-grained rationales, Shridhar et al. (2023) introduce Socratic CoT by decomposing the original question into a series of subquestion-solution pairs and generating CoT for them separately. Additionally, Kang et al. (2024) propose a neural reranker to acquire supplementary relevant documents for rationale generation in knowledge-intensive reasoning tasks.
Human-like Rationale. Another intriguing avenue in synthesized rationale delves into making the reasoning process more human-like. Many studies emulate human diverse thinking in problem-solving, sampling multiple reasoning pathways for a given question Gao et al. (2021); Wang et al. (2022b); Chen et al. (2023f); Liu et al. (2023c). Subsequent studies Tong et al. (2023); Balepur et al. (2023); Ma and Du (2023) explore the elimination reasoning in LLMs, checking each reasoning pathway reversely and removing the incorrect candidates. Moreover, various works Yin et al. (2023); Liang et al. (2023); Xu et al. (2023d); Liu et al. (2023e) explore the peer collaboration and debate among individual LLMs to capture human-like discussions as rationales.
@@ 3.3 Pairwise Feedback
Ranking with LLMs. One technique is to sample multiple responses and have the LLM rank these candidates based on various criteria Bai et al. (2022); Lee et al. (2023b); Yuan et al. (2024). Sun et al. (2024a) sample two responses from the initial policy model and use the model to select the preferred response based on a human-written principle Sun et al. (2024b). Zhang et al. (2024a) propose a self-evaluation mechanism, generating questions for each response and measuring factuality by the LLM’s confidence in the answers. To improve synthetic data quality, Pace et al. (2024) combine the Best-of-N and Worst-of-N sampling strategies and introduce the West-of-N approach. They constructed data pairs by identifying the best- and worst-scored responses according to a pre-trained preference model. In robotics, Zeng et al. (2024) iteratively update the reward function with the self-ranked responses from LLMs, enhancing learning efficiency without human supervision.
@@ 3.5 Other Domain-specific Data
Distilling multi-round conversations from LLMs presents a highly cost-effective approach for constructing high-quality dialogue datasets Kim et al. (2023a); Xu et al. (2023b); Chen et al. (2023b); Li et al. (2024d) or enhancing existing ones Zheng et al. (2023); Chen et al. (2022); Zhou et al. (2022a). In graph and tabular data, several studies prompt LLMs to contextualize these structural data Xiang et al. (2022); Kim et al. (2023a); Li et al. (2024b); Ronzano and Nanavati (2024) or distill structural insights from raw text Bi et al. (2024); Li et al. (2024c); Ding et al. (2024); Xiong et al. (2024); Tuozzo (2022). Moreover, LLMs have also been widely adopted in the research of robotics and agents, serving as proficient data annotators to generate plans Huang et al. (2022); Brohan et al. (2023); Rana et al. (2023); Singh et al. (2023); Lin et al. (2023a), simulation tasks Wang et al. (2023a); Ha et al. (2023) and supervised signal Kwon et al. (2022); Du et al. (2023b); Tan et al. (2024b, a). Besides, LLMs are acting as efficient data annotators in various artificial intelligence domains, including multi-modal Li et al. (2023f); Yin et al. (2024); Chen et al. (2024a); Li et al. (2024f), recommendation system Acharya et al. (2023); Shen et al. (2024); Wei et al. (2024); Zhang et al. (2024b), information extraction Josifoski et al. (2023); Jeronymo et al. (2023); Li et al. (2024a); Ma et al. (2024); Bonn et al. (2024) and etc Chu et al. (2024a); Bhattacharjee et al. (2024); Martorana et al. (2024); Hsu et al. (2024).

### v3

# Large Language Models for Data Annotation and Synthesis: A Survey

@@ Abstract
Data annotation and synthesis generally refers to the labeling or generating of raw data with relevant information, which could be used for improving the efficacy of machine learning models. The process, however, is labor-intensive and costly. The emergence of advanced Large Language Models (LLMs), exemplified by GPT-4, presents an unprecedented opportunity to automate the complicated process of data annotation and synthesis. While existing surveys have extensively covered LLM architecture, training, and general applications, we uniquely focus on their specific utility for data annotation. This survey contributes to three core aspects: LLM-Based Annotation Generation, LLM-Generated Annotations Assessment, and LLM-Generated Annotations Utilization. Furthermore, this survey includes an in-depth taxonomy of data types that LLMs can annotate, a comprehensive review of learning strategies for models utilizing LLM-generated annotations, and a detailed discussion of the primary challenges and limitations associated with using LLMs for data annotation and synthesis. Serving as a key guide, this survey aims to assist researchers and practitioners in exploring the potential of the latest LLMs for data annotation, thereby fostering future advancements in this critical field [^1] We release the paper list of LLM-based data annotation and synthesis at https://github.com/Zhen-Tan-dmml/LLM4Annotation.
@@ 1 Introduction
In the complex realm of machine learning and natural language processing (NLP), data annotation and synthesis stand out as a critical yet challenging task, extending beyond simple label attachment to encompass a diverse array of fundamental or auxiliary information. This detailed process typically involves ❶ categorizing raw data with class or task labels for basic classification, ❷ adding intermediate labels for contextual depth Yu et al. (2022), ❸ assigning confidence scores to assess annotation reliability Lin et al. (2022), ❹ applying alignment or preference labels to tailor outputs to specific criteria or user needs, ❺ annotating entity relationships to understand how entities within a dataset interact with each other Wadhwa et al. (2023), ❻ marking semantic roles to define the underlying roles that entities play in a sentence Larionov et al. (2019), ❼ tagging temporal sequences to capture the order of events or actions Yu et al. (2023), or ❽ Synthesize data in the format of instruction Wang et al. (2022b), response Zhang and Yang (2023a), reasoning Wang et al. (2022a), pairwise Bai et al. (2022) and textual feedback Pan et al. (2024) to for language model tuning.
Despite its wide applications, data annotation and synthesis poses significant challenges for current machine learning models due to the complexity, subjectivity, and diversity of data Yang et al. (2023d). This process requires domain expertise and is resource-intensive, particularly when manually labeling or creating large datasets. Advanced LLMs such as GPT-4 OpenAI (2023), Gemini Team et al. (2023), and LLaMA-2 Touvron et al. (2023b) offer a promising opportunity to revolutionize data annotation. LLMs serve as more than just tools but play a crucial role in improving the effectiveness and precision of data annotation. Their ability to automate annotation tasks, ensure consistency across large volumes of data (Hou et al., 2023), and adapt through fine-tuning or prompting for specific domains (Song et al., 2023; Zhang et al., 2024a), significantly mitigates the challenges encountered with traditional annotation and synthesis methods, setting a new standard for what is achievable in the realm of NLP. This survey delves into the nuances of using LLMs for data annotation and synthesis, exploring methodologies, utilizing strategies, and associated challenges in this transformative approach. Through this exploration, we aim to shed light on the motivations behind embracing LLMs as catalysts for redefining the landscape of data annotation and synthesis in machine learning and NLP. We explore the utilization of LLMs for annotation synthesis in this survey, making four main contributions:
  • LLM-Based Annotation Generation: We dive into the process of synthesizing annotations for various data types, including instruction & response, rationale, pairwise feedback, textual feedback, and other domain-specific data. Additionally, we discuss the criteria (e.g., diversity and quality) in the annotation process.
@@
Differences from Other LLM-related Surveys. While existing surveys in the NLP domain extensively cover architectural nuances Zhao et al. (2023a), training methodologies Liu et al. (2023d), and evaluation protocols Chang et al. (2023) associated with LLMs, their main focus lies on the capabilities of models for specific end tasks such as machine translation Min et al. (2021), alignment Wang et al. (2023g), code generation Zan et al. (2023), and medical analysis Thirunavukarasu et al. (2023). In contrast, this survey distinguishes itself by focusing primarily on the application of these potent next-generation LLMs to the intricate realm of annotation synthesis, a domain that is crucial yet underexplored.
@@ 2 Preliminaries
In this section, we delve into our approach to the annotation synthesis process. We introduce two core models: an annotator model, denoted as $\mathcal{A}$, which maps input data to annotations, and a task learner, represented as $\mathcal{L}$, that utilizes or learns from these annotated data to accomplish specific tasks. Our primary focus is on utilizing advanced LLMs like GPT-4 (OpenAI, 2023) and LLaMA (Touvron et al., 2023a) as annotators ($\mathcal{A}$), while the task learner ($\mathcal{L}$) can be another large model Chiang et al. (2023a) or a less complex one such as BERT (Devlin et al., 2018), which utilizes these annotated data to perform designated tasks. LLM-generated annotations encompass categorical labels and enhance raw data points with a comprehensive array of auxiliary signals. These annotations, including confidence scores, contextual details, and other metadata, extend beyond traditional categorical labels.
@@ 3 LLM-Based Annotation Generation
The emergence of LLMs has sparked significant interest in their capacity for high-quality, context-sensitive annotation synthesis. This section discusses various kinds of annotations and data produced via LLMs.
@@ 3.1 Instruction & Response
Instruction and response are the two fundamental components that constitute a dataset for LLM fine-tuning and in-context learning (ICL). Previous NLP datasets Li et al. (2017); Wang et al. (2018); Ouyang et al. (2022) mainly rely on human annotators to construct. Recently, with the advent of LLMs, automatic and generative methods Meng et al. (2022); Ye et al. (2022a, b); Wang et al. (2024f); Wu et al. (2024b); Liu et al. (2024a) have gained more focus in data annotation.
@@
Instruction Diversity. The diversity of instruction has been proven crucial for LLM learning Li et al. (2023e); Song et al. (2024b, a); Tang et al. . Recent studies have explored various methods to diversify and augment instructions in the original datasets. For example, Yoo et al. (2021) enhance data diversity by mixing two different samples to create a new one. Wang et al. (2022b) use a few manually-written seed instructions and iteratively augment them with a generate-then-filter pipeline. Additionally, Meng et al. (2023); Wang et al. (2023f) train an instruction generation model in the original dataset to augment the diversity of instruction. Gupta et al. (2023) employ a multi-step prompting method to first generate task descriptions, which are then used as instance seeds to guide LLMs in instruction generation. To obtain informative and diverse examples, Wang et al. (2023c) propose an explain-then-generate pipeline with LLMs for iterative data synthesis. Besides, Li et al. (2023a) paraphrase the given sample multiple times to help LLMs understand them from different perspectives. Köksal et al. suggest a clustering-based data selection method to ensure diversity in the initial seed data for augmentation. Recently, Yu et al. (2024) introduce AttrPrompt as an effective way to balance diversity and cost in LLM-based data annotation. Xu et al. (2024) propose to synthesize high-quality instruction data at scale by extracting it directly from an aligned LLM and present a self-synthesis method for generating large-scale alignment data named Magpie. To improve the diversity, Chan et al. (2024) introduce Persona Hub – a collection of 1 billion diverse personas automatically curated from web data, to foster the creation of diverse synthetic data at scale for various scenarios. Zhu et al. (2024) introduce FANNO, a fully autonomous, open-sourced framework that revolutionizes the annotation process without the need for pre-existing annotated data. Recently, Kowshik et al. (2024) propose CorrSynth, which generates data that is more diverse and faithful to the input prompt using a correlated sampling strategy. To augment the clinical data safely and protect the patient’s privacy,  propose to use GPT-4 for data augmentation through one-shot and zero-shot prompts.  Divekar and Durrett (2024) propose Synthesize by Retrieval and Refinement (SynthesizRR), which uses retrieval augmentation to introduce variety into the dataset synthesis process: as retrieved passages vary, the LLM is seeded with different content to generate its examples.  Li et al. (2024a) introduce a general and scalable framework, IDEA-MCTS (Instruction Data Enhancement using Monte Carlo Tree Search), a scalable framework for efficiently synthesizing instructions.
@@
Response Quality. High-quality responses are essential for effective fine-tuning and ICL Luo et al. (2024a). To improve the quality of the generated response, Zhang and Yang (2023a) frame the response generation as reading comprehension tasks and create detailed prompts for LLMs. Huang et al. (2023) adopt self-consistency Wang et al. (2022b) in response generation, selecting from the candidate response with the highest confidence score. Furthermore, Yang et al. (2024b) propose self-distill and augment the instruction tuning dataset by rewriting the original responses. Pang et al. (2024b) conduct social simulations to ensure high-quality, human-valued responses from LLMs. Moreover, Liu et al. (2024c) introduce a multi-step prompting including question analysis, answer guidance and safe answer production in their response generation pipeline. Guo et al. (2024a) enhance the LLMs outputs’ quality by implementing retrieval-augmented ICL and providing LLMs with relevant documents. To ensure LLMs provide responses aligned with human values, Sun et al. (2024b) and Wang et al. (2024a) conduct principle-driven prompting, guiding LLMs with well-crafted and detailed principles. Besides, Lupidi et al. (2024) propose Source2Synth, which takes as input a custom data source and produces synthetic data points with intermediate reasoning steps grounded in real-world sources. Recently, Chen et al. (2024a) follow “less is more” principle, achieving better results with fewer data points in mathematical reasoning data generation.
3.2 Label
Label is an important component of the traditional classification task in NLP. Nowadays, many researchers focus on automating label annotation with the assistance of LLMs Yadav et al. (2024); Tseng et al. (2024). Chen et al. (2024b) introduce an innovative approach where we employ LLMs as expert annotators for event extraction. Martorana et al. (2024b) propose a method to support metadata enrichment using topic annotations generated by several LLMs. Both Wu et al. (2024a) and Ahmed et al. (2024) explore the potential of large language models (LLMs) as automated data annotators to improve efficiency and consistency in label annotation tasks. One interesting work from Li et al. (2023b) proposes CoAnnotating, a novel paradigm for Human-LLM co-annotation of unstructured texts at scale. Moreover, Tekumalla and Banda (2023) evaluate the utilization of LLM in labeling COVID-19 vaccine-related tweets, with the purpose of comparing performance against human annotators. To address the potential limitation of LLMs’ annotation, Törnberg (2024) propose a comprehensive set of standards and best practices for their reliable, reproducible, and ethical use. Additionally, there are also some works that utilize LLMs to improve the original annotation made by human annotators Laskar et al. (2023); Flamholz et al. (2024); Wang et al. (2024e). To reduce costs, Schmidt et al. (2024) argue that domain-agnostic knowledge from LMs, such as linguistic understanding, is sufficient to create a well-curated dataset. Recently, Choi et al. (2024a) propose a novel approach to universal domain generalization that generates a dataset regardless of the target domain.  Li et al. (2024h) integrate LLMs with a novel sorting method to address the multi-level function call relationships within repositories for code annotation.  Choi et al. (2024b) leverage approaches such as chain-of-thought and majority voting to imitate human annotation and classify unrelated documents from the Multi-News dataset, which is widely used for the multi-document summarization task. To calibrate models and improve their generalization, Ba et al. (2024) use LLMs as synthetic data generation strategies to lower the ECE bound and improve model accuracy on real test data.
@@ 3.3 Rationale
Rationale Structure. Following Kojima et al. (2022), there is a notable interest in abstracting the reasoning process of LLMs into diverse structures and format, including trees Hao et al. (2023); Yao et al. (2024), graphs Besta et al. (2024); Yao et al. (2023), tables Wang et al. (2024g), programs Chen et al. (2023e), recursion Qi et al. (2023), and concepts Tan et al. (2023).
Rationale Quality. To produce high-quality and fine-grained rationale, diverse methodologies have been employed. Wang et al. (2022a) prompt frozen LLMs to produce choice-specific rationales to elucidate each choice in a sample. Wang et al. (2023b) employ contrastive decoding to foster more plausible rationales, taking into account gold-standard answers. Liu et al. (2023a) curate meticulously designed prompts to derive high-quality rationales from GPT-4 and construct a logical CoT instruction tuning dataset. For attaining fine-grained rationales, Shridhar et al. (2023) introduce Socratic CoT by decomposing the original question into a series of subquestion-solution pairs and generating CoT for them separately. Additionally, Kang et al. (2024) propose a neural reranker to acquire supplementary relevant documents for rationale generation in knowledge-intensive reasoning tasks. Besides, Zhou et al. (2024) explore the potential and limitations of using graph-based synthetic reasoning data as training signals to enhance LLMs’ reasoning capabilities. Recently, Wei et al. (2024b) propose InstructRAG, where LMs explicitly learn the denoising process through self-synthesized rationales.
Human-like Rationale. Another intriguing avenue in synthesized rationale delves into making the reasoning process more human-like Gao et al. (2023). Many studies emulate human diverse thinking in problem-solving, sampling multiple reasoning pathways for a given question Gao et al. (2021); Wang et al. (2022b); Chen et al. (2023f); Liu et al. (2023c). Subsequent studies Tong et al. (2023); Balepur et al. (2023); Ma and Du (2023) explore the elimination reasoning in LLMs, checking each reasoning pathway reversely and removing the incorrect candidates. Moreover, various works Yin et al. (2023); Liang et al. (2023); Xu et al. (2023d); Liu et al. (2023e) explore the peer collaboration and debate among individual LLMs to capture human-like discussions as rationales.
@@ 3.4 Pairwise Feedback
Ranking with LLMs. One technique is to sample multiple responses and have the LLM rank these candidates based on various criteria Bai et al. (2022); Lee et al. (2023b); Yuan et al. (2024). Sun et al. (2023b) sample two responses from the initial policy model and use the model to select the preferred response based on a human-written principle Sun et al. (2024b). Zhang et al. (2024b) propose a self-evaluation mechanism, generating questions for each response and measuring factuality by the LLM’s confidence in the answers. To improve synthetic data quality, Pace et al. (2024) combine the Best-of-N and Worst-of-N sampling strategies and introduce the West-of-N approach. They constructed data pairs by identifying the best- and worst-scored responses according to a pre-trained preference model. In robotics, Zeng et al. (2024) iteratively update the reward function with the self-ranked responses from LLMs, enhancing learning efficiency without human supervision.
@@ 3.6 Other Domain-Specific Data
Distilling multi-round conversations from LLMs presents a highly cost-effective approach for constructing high-quality dialogue datasets Kim et al. (2023a); Xu et al. (2023b); Chen et al. (2023b); Li et al. (2024f); Wang et al. (2024d); Liang et al. (2024a) or enhancing existing ones Zheng et al. (2023a); Chen et al. (2022); Zhou et al. (2022a); Sun et al. (2024a). In graph and tabular data, several studies prompt LLMs to contextualize these structural data Xiang et al. (2022); Kim et al. (2023a); Li et al. (2024d); Ronzano and Nanavati (2024); Xiong et al. (2023b, 2024b) or distill structural insights from raw text Bi et al. (2024); Li et al. (2024e); Ding et al. (2024); Xiong et al. (2024a); Tuozzo (2022). Moreover, LLMs have also been widely adopted in the research of robotics and agents, serving as proficient data annotators to generate plans Huang et al. (2022); Brohan et al. (2023); Rana et al. (2023); Singh et al. (2023); Lin et al. (2023a), simulation tasks Wang et al. (2023a); Ha et al. (2023) and supervised signal Kwon et al. (2022); Du et al. (2023b). Besides, LLMs are acting as efficient data annotators in various artificial intelligence domains, including multi-modal Li et al. (2023f); Yin et al. (2024); Chen et al. (2024c); Luo et al. (2024b); Liu et al. (2024b); Wang et al. (2024c); Cheng et al. (2024a), recommendation system Acharya et al. (2023); Shen et al. (2024); Wei et al. (2024a); Zhang et al. (2024c), information extraction Josifoski et al. (2023); Jeronymo et al. (2023); Li et al. (2024b); Ma et al. (2024); Bonn et al. (2024), multilingual annotation Frei and Kramer (2023); Hamerlik et al. (2024); Latouche et al. (2024) and etc Chu et al. (2024a); Bhattacharjee et al. (2024); Martorana et al. (2024a); Zhao et al. .